Phillip Burton's savings account has a balance of 1,531. After 3 years what will the amount of interest be at 3% compound quarterly?
Answer by stanbon(75887)

Phillip Burton's savings account has a balance of 1,531. After 3 years what will the amount of interest be at 3% compound quarterly?
-----------------
I = P(1+(r/n))^(nt) - P
-----
Interest = 1531[(1+(0.03/4))^(4^3))-1] = 1531*0.09381 = $143.62
-------
Cheers,
Stan H.
---------------